11. If a metal surface is irradiated with 185 nm light, electrons are ejected with a velocity of 8.90 x 105 m/s.

What is the minimum frequency (in Hz) of light that could be used which will still eject electrons from the surface?
Answer #1

11)

Given:

lambda = 1.85*10^-7 m

use:

E = h*c/lambda

=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(1.85*10^-7 m)

= 1.074*10^-18 J

KE = 0.5*m*v^2

= 0.5*(9.11*10^-31 Kg)*(8.90*10^5 m/s)^2

= 3.61*10^-19 J

Now use:

KE = E - Eo

3.61*10^-19 J = 1.074*10^-18 J - Eo

Eo = 7.13*10^-19 J

use:

Eo = h*fo

7.13*10^-19 J =(6.626*10^-34 J.s)*fo

fo = 1.076*10^15 Hz

Answer: 1.08*10^15 Hz